A pendulum has a mass of 2.35 kg. If the pendulum swings back and forth 110 times in 90.0 seconds, what is the length of the pendulum


1
Expert's answer
2022-01-13T09:20:52-0500

"T=2\pi\sqrt{\frac lg},"

"T=\frac tn,"

"l=\frac{gt^2}{4\pi^2n^2}=16.6~cm."